OpenStudy (anonymous):

Show all work in dividing 9 x squared minus 1 all over 8 x minus 4 divided by the fraction 3 x squared plus 5 x minus 2 all over 2 x squared plus x minus 1 (list restrictions).

OpenStudy (anonymous):

\[\frac{ 9x^2-1 }{8x-4 } \div \frac{3x^2+5x-2 }{2x^2+x-1 }\] I know that since this is a division problem, you'd use the reciprocal of the second fraction and it would now look like: \[\frac{ 9x^2-1 }{ 8x-4 } \div \frac{ 2x^2+x-1 }{ 3x^2+5x-2 }\] I need help from here. Thanks :)

OpenStudy (anonymous):

Sorry that should be a multiplication symbol in the second problem.

undeadknight26 (undeadknight26):

o.o Omg imnot very good with these -_- @shrutipande9 @wolfe8 @thomaster

OpenStudy (wolfe8):

Recall that \[\frac{ a }{ b }\div \frac{ c }{ d }=\frac{ a }{ b }\times \frac{ d }{ c }\]

OpenStudy (wolfe8):

And then you factorize what you can and see if you can cancel out anything. At the end, multiply top with top and bottom with bottom.

OpenStudy (anonymous):

For the first part I have: \[\frac{ (3x+1)(3x-1) }{ 4(2x-1) } \] How do you factor out the second fraction?

OpenStudy (anonymous):

@wolfe8

OpenStudy (wolfe8):

I was working it out. I got \[\frac{ (x+1)(2x-1) }{ (x+2)(3x-1) }\]

OpenStudy (anonymous):

is that for the second fraction?

OpenStudy (wolfe8):

Yeah

OpenStudy (anonymous):

So now I multiply right?

OpenStudy (wolfe8):

First you can cancel out what appears on the inverse side that are the same. Get what I mean? like if you have a/b*c/a you can cancel out the a's to have 1/b*c

OpenStudy (anonymous):

so I'm left with \[\frac{ (3x+1) }{4 } \times \frac{ (x+1) }{ (x+2)}\] Now do I multiply?

OpenStudy (wolfe8):

Yup looks like it!

OpenStudy (anonymous):

So I got: \[\frac{ 3x^2+3x+1 }{ 4(x+2) }\] is that right?

OpenStudy (wolfe8):

I think it's supposed to be 4x on top not 3x.

OpenStudy (anonymous):

Oh you're right, I found my mistake

OpenStudy (wolfe8):

Yup. Other than that I think you got it.

OpenStudy (anonymous):

I don't have to factor the numerator do I?

OpenStudy (wolfe8):

What do you mean by factor? You can't factor it any further but you can remove the brackets. I think you should too.

OpenStudy (anonymous):

I meant like in the final answer the 3x^2+4x+1 I was asking if I had to factor that

OpenStudy (wolfe8):

You don't have to. If you want to, you just had to leave them in the brackets when in the final multiplication step.

OpenStudy (anonymous):

Oh alright. Thank you so much! :)

OpenStudy (wolfe8):

You're welcome. Good job.